OpenStudy (anonymous):

. The Bunker Hill Monument in Charlestown' Massachusetts, is 22l feet tall. It casts a shadow 189 feet long at the same time that a nearby tree casts a shadow 29 feet long' How tall is the tree to the nearest foot?

OpenStudy (opcode):

Proportions. \(\huge\frac{221}{189} \times \frac {x}{29}\) \(\huge 189x = 6409\) Divide each side by 189.

OpenStudy (anonymous):

Do I solve?

OpenStudy (opcode):

\(\huge 6409 / 189 = 33.9100529 \) Round 33.9100529 to the nearest foot.

OpenStudy (anonymous):

oh. ok

OpenStudy (anonymous):

one second.

OpenStudy (anonymous):

34 ft?

OpenStudy (opcode):

Yes sir.

OpenStudy (anonymous):

Lol. thanks kami